With Binomial Distributions and Probability:
In a lottery, each ticket buyer has a 1 in 10 chance of winning a prize. If Steve buys 10 tickets, what is the probability that he will win at least 1 prize? How do you work this problem out?

Answer by stanbon(57984)

With Binomial Distributions and Probability:
In a lottery, each ticket buyer has a 1 in 10 chance of winning a prize. If Steve buys 10 tickets, what is the probability that he will win at least 1 prize? How do you work this problem out?
-------------------------------
Binomial Problem:
n = 10 ; p = 1/10
---
P(win at least 1) = 1 - P(win none)
----
P(win at least 1) = 1 - (0.9)^10
---
P(win at least 1) = 1 - 0.3487 = 0.6513
========================================
Cheers,
Stan H.